Previous

Calendar

October

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

28

JD272

29

JD273

30

JD274
Site Visit (JB, NT, SpB): Details , Photos (28)

span @ 1450, N2 @ 1200

swapped 15 flasks at west robot

added 2 accelerometers to the "fast" logger

started running ops47 (added 2 accelerometers)

started digging holes for the CS616 probes

1

JD275
Site Visit (JB, NT, SpB): Details , Photos (12)

tested the replacement laptop at the tower

tested some CS616 probes with the soil logger (see details)

2

JD276

3

JD277
Site Visit (JB, SpB, MP, NT, DB, DR-I, SO): Details , Photos (11)

JB, SpB and MP finished removing the PAR wands from grove 1 and started work on grove 2, hauled the adam and some other par equipment back to the trailer

SpB troubleshot the CS616s some more...

4

JD278

5

JD279

6

JD280

7

JD281
Site Visit (JB, NT, SpB, MP, AW, LL): Details , Photos (22)

swapped 15 flasks at the west robot

finished par wand removal at grove 2

SpB got five of the CS616s working and calibrating in some soil near the soil logger

8

JD282
Swap Teal Jeep for Gray Jeep

9

JD283
Site Visit (JB, NT, SpB, MP, AW): Details , Photos (28)

finished par wand removal at grove 3

connected all 8 new CS-616s to soil logger 416 mux, started IC (see notes for details)

from 12:15-12:30 MST, removed 5.6m csat (sn 0328) from mini-tower

Sunny and clear at the site (upslope winds). Socked in with clouds down on the plains.

10

JD284

mailed 5.6m csat (s/n 0328) for re-cal and rain upgrade

at 20:00 MST, N2 calibrations with LI-6262 (6251 too?) system started having problems.

11

JD285
First Significant snow of the season (approx 3-4").

12

JD286
at times on the edge of colder air from the lowlands (approx 9AM MST).

13

JD287

14

JD288
Site Visit (JB, NT): Details

brought up 1 new N2 cylinder (CC199722) and 1 new span cylinder (CC282360) to the trailer from Ramaley

swapped 15 flasks at the west robot then changed it to "winter mode" (i.e. 4 flasks every two weeks)

downloaded from soil, UC, UCB, nctc and prof

fired up the mrs logger and Noah's logger, sent mrs_03b.dld to the mrs logger

span @ 1250 psi, N2 @ 700 psi

see "details" about the solenoid valves in the LI-6262 box

from ~1:55-1:58 MDT did a manual N2 calibration of the EC system to test the N2 flow rate

15

JD289
AmeriFlux Meeting (in Boulder).

16

JD290
AmeriFlux Meeting. Site Visit (JB, SpB): Details

brought up the last two cylinders from Ramaley, both span, CC153480 and CC245057

N2 (CC125575) bled almost dry since Tuesday! a leak somewhere in the system...retired this cylinder

started new N2 cylinder CC262287 at 1850 psi

span @ 1250 psi

swapped the mag. perch. and indicator tube at the LI-6251

disconnected the N2 zero cal at the LI-6262, left everything else as-is (to try and isolate any leak).

17

JD291

ordered three tanks of N2 from Airgas

AmeriFlux Meeting. Gray Jeep back to motor pool

18

JD292

19

JD293
Site Visit (JB, SpB, MP, NT, SL, Flurry, Harley): Details , Photos (22)

Installed the new soil moisture sensors (cs-616s) and re-deployed some of the cs-615 sensors in new locations so that each one is co-located with a snow depth sensor (see details)

N2 cylinder was at approx 1830 psi (so leak is most likely in the N2 zero cal of the LI-6262).

made an amp connector extension with 75 ft. of cable for one of the new snow temp sensors, but the sensor didn't seem to work...maybe not the same one we tested at the mrs logger?

20

JD294

21

JD295

22

JD296

23

JD297

24

JD298
Site Visit (JB, SpB): Details , Photos (20)

at 11:40 MDT, started ops48 (added extra co2 param and removed one accelerometer)

N2 @ 1700, span @ 1100

got a snow temp probe working with mrs_03c.dld (see details for wiring)

25

JD299

26

JD300

27

JD301

28

JD302
Site Visit (JB, SpB, MP, NT): Details

N2 @ 1610, span @ 1150

cleaned the krypton

see details for additional notes

swapped 4 flasks at west robot

29

JD303

30

JD304
Site Visit (JB, MP, NT): Details

did all of Maggie's new growth collections

helped winterize NT's plots

checked on the two tanks running at the tower, both okay

brought up a new N2 tank from Ramaley, cc264964

31

JD305
Site Visit (JB, SpB): Details , Photos (32)

brought up a new N2 tank from Ramaley, cc264970

from 10:30-11:15 MST, removed csat 0438 (ver3) at 21.5m, added csat 0328 at 21.5m (w/ ver4, "rain" upgrade)

at 14:30 MST, added csat 0438 at 5.7m

at 14:30 MST, started ops49 (added a TC at 30cm AGL to UCB logger)

at 14:30 MST, started running "ucb_10.csi" on UCB logger

Halloween

1

JD306